OpenStudy (anonymous):

A proportional relationship between the number of pounds of cabbage (x) and the price in dollars (y) is graphed, and the ordered pair (5, 3) is on the graphed line. Part A: What is the price of 1 pound of cabbage? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (10, 6) on the graph represent? Explain in words.

OpenStudy (anonymous):

please can you hurry up

OpenStudy (texaschic101):

(5,3)...x = 5 and y = 3 proportions are nothing but equivalent fractions. pounds of cabbage 5 ---------------- = ---- cost of cabbage 3 A. 5/3 = 1/x --- 5 lbs of cabbage to $3 = 1 lb of cabbage to $x this is your proportion....now we cross multiply to solve for x. (5)(x) = (3)(1) 5x = 3 x = 3/5 x = 0.6 so 1 lb of cabbage costs 0.60 cents B. (10,6) --- this means 10 lbs of cabbage costs $6

OpenStudy (anonymous):

please can you hurry up

OpenStudy (anonymous):

please can you hurry up

OpenStudy (anonymous):

please can you hurry up

OpenStudy (anonymous):

please can you hurry up

OpenStudy (anonymous):

please can you hurry up

OpenStudy (anonymous):

please can you hurry up